# A bi-fidelity method for the uncertain Vlasov-Poisson system near quasineutrality in an asymptotic-preserving PIC framework

### v1

@@ 4 A bi-fidelity method for model with random parameters
In this section, we briefly review the bi-fidelity method. Assume the expensive high-fidelity model and the cheap low-fidelity model are available to generate the high-fidelity solution $\bm{U}^{H}(z)$ and the low-fidelity solution $\bm{U}^{L}(z)$ respectively, for any given parameter $z$. We refer to [40, 49] for more details and summarize the Algorithm 1 as below:
Offline :
@@
1Select a sample set $\Gamma=\{z_{1},z_{2},\ldots,z_{M}\}\subset I_{z}$.
2Run the low-fidelity model $\bm{U}^{L}(z_{j})$ for each $z_{j}\in\Gamma$ ($j=1,\cdots,M$).
@@
3Select $r$ “important” points from $\Gamma$ and denote it by $\gamma=\{z_{1},\cdots z_{r}\}\subset\Gamma$. Construct the low-fidelity approximation space $\mathscr{U}^{L}(\gamma)$.
@@
4Run high-fidelity simulation at each point in the selected sample set $\gamma$. Construct the high-fidelity approximation space $\mathscr{U}^{H}(\gamma)$.
Online :
For any given $z$, compute the low-fidelity solution $U^{L}(z)$ and the corresponding low-fidelity coefficients by projection:
$\bm{U}^{L}(z)\approx\mathcal{P}_{\mathscr{U}^{L}\left(\gamma\right)}\left[\bm{U}^{L}(z)\right]=\sum_{k=1}^{r}c_{k}^{L}(z)\bm{U}^{L}\left(z_{k}\right).$ (4.1)
Construct the bi-fidelity approximation by applying the same approximation rule as in low-fidelity model:
$\bm{U}^{B}(z)=\sum_{k=1}^{r}c_{k}^{L}(z)\bm{U}^{H}(z_{k}).$ (4.2)
Algorithm 1 A bi-fidelity approximation
@@
Our sought bi-fidelity approximation of $u^{H}$ is constructed by the equation (4.2). In the offline stage, one employs the cheap low-fidelity model to select the most important parameter points $\gamma=\{z_{1},\cdots,z_{r}\}$ by the greedy procedure [16]. In the online stage, for any sample point $z\in I_{z}$, we project the low-fidelity solution $\bm{U}^{L}(z)$ onto the low-fidelity approximation space $\mathscr{U}^{L}(\gamma)$ by using (4.1), where $\mathcal{P}_{\mathscr{U}^{L}(\gamma)}$ is the projection operator onto $\mathscr{U}^{L}(\gamma)$ with projection coefficients $c_{k}$ computed by the Galerkin approach:
${\bf G}^{L}{\bf c}={\bf f},\,\text{\leavevmode\nobreak\ with\leavevmode\nobreak\ }{\bf f}=(f_{k})_{1\leq k\leq r}\,\text{\leavevmode\nobreak\ and\leavevmode\nobreak\ }f_{k}=\langle\bm{U}^{L}(z),\bm{U}^{L}(z_{k})\rangle^{L},$
where ${\bf G}^{L}$ is the Gramian matrix for $\bm{U}^{L}(\gamma)$.
High-fidelty solver.
@@
We desire to solve the VPME system with multi-dimensional uncertain parameters, thus in our bi-fidelity method, the VPME system (2.1c) is considered as the high-fidelity model. We employ the AP-PIC method studied in subsection 3.3 as our high-fidelity solver.
Low-fidelity solver.
@@
We choose the macroscopic Euler system (2.4) that characterizes a simpler physics as our low-fidelity model in the bi-fidelity method. We use the finite difference method (FDM) to discretize the spatial space and forward Euler scheme for the time discretization. Consider the conservation form of (2.4) for one-dimensional space and velocity:
$\frac{\partial\bm{U}}{\partial t}+\frac{\partial\bm{F}(\bm{U})}{\partial x}=0,\quad\bm{U}=\begin{pmatrix}n\\ nu\end{pmatrix},\quad\bm{F}(\bm{U})=\begin{pmatrix}nu\\ nu^{2}+n\end{pmatrix}.$ (4.3)
@@
where $n=n(x,t),\>u=u(x,t)$ are the ion density and velocity given in (2.3) in the computational domain $\left[x_{L},x_{R}\right]$. Let $h$ be a uniform space step and $\tau$ be a uniform time step, then $x_{j}=jh,\>j=0,1,2,\dots,N_{l}$ with $x_{L}=x_{0},\>x_{R}=x_{N_{l}}$ and $t^{n}=n\tau,\>n=1,2,\dots,M$.
@@
In our flux construction, Lax-Friedrichs scheme–as one of the most popular finite difference methods for solving hyperbolic partial differential equations [35, 46, 47] is used. Let $\bm{U}^{n}_{j}=\bm{U}(x_{j},t^{n})$ be the approximated values of $\bm{U}$ at position $x_{j}$ and time $t^{n}$. The time discretization for (4.3) becomes
$\bm{U}^{n+1}_{j}=\bm{U}^{n}_{j}-\frac{\tau}{h}(\bm{F}^{n}_{j+\frac{1}{2}}-\bm{F}^{n}_{j-\frac{1}{2}}),$ (4.4)
where $\bm{F}^{n}_{j+\frac{1}{2}}$ is given by
$\bm{F}^{n}_{j+\frac{1}{2}}=\frac{1}{2}\left(\bm{F}(\bm{U}^{n}_{j+1}\right)+\bm{F}(\bm{U}^{n}_{j}))-\frac{h}{2\tau}(\bm{U}^{n}_{j+1}-\bm{U}^{n}_{j}).$ (4.5)
@@
$\bm{U}^{n+1}_{j}=\frac{\bm{U}^{n}_{j+1}+\bm{U}^{n}_{j-1}}{2}-\frac{\tau}{2h}(\bm{F}(\bm{U}^{n}_{j+1})-\bm{F}(\bm{U}^{n}_{j-1})).$ (4.6)
@@
5 Numerical examples
In our numerical experiments, we first study the deterministic VPME system, then extend to the problems with random uncertainties and validate the efficiency of our bi-fidelity method. For all the following experiments, the regularization parameter $\eta$ in (3.1d) and the error tolerance $\eta_{1}$ in the Newton’s method (3.30) for solving the reformulated Poisson are set as $\eta=1,\,\eta_{1}=1\times 10^{-8}$. For notational convenience, we name the reformulated VPME system (3.15)–(3.16) as the RF-VPME system, and its quasineutral limit (3.17)–(3.18) as the Quasi-VP system.
5.1 The RF-VPME system: an explicit scheme
@@
Set the final time as $t=2$ and $\varepsilon=0.01$ in the model. Let $v\in[-4,4]$ with uniform mesh and the number of grid points be $N_{v}=800$, the spatial mesh size be $N_{x}=200$ and time step be $\Delta t=0.2\Delta x$. The density $n$ and velocity $u$ are shown in Fig. 2, where the blue line represents solution of the explicit scheme (3.36) based on building a penalty for the nonlinear term $e^{\phi^{m+1}}$, and the red line stands for the implicit scheme using Newton’s iteration (3.29). The solutions of both methods match well, and both schemes are stable for long time. In our future work, we will incorporate this explicit scheme with the PIC method for solving the Vlasov equation (2.1a) and develop higher-order AP schemes.
Figure 2: (Two schemes for solving the reformulated Poisson, $\varepsilon=0.01$). Density $n$ (left) and velocity $u$ (right) at $t=2$ by an implicit scheme (3.23) with Newton’s iteration (red) and an explicit scheme based on penalty (3.36) (blue).
5.2 The RF-VPME system: deterministic problem
@@
We particularly study the cases when $\varepsilon=1,0.1,0.01$ and $10^{-4}$. The final time is $t=0.2$. When $\varepsilon=1$, the numerical solution is obtained by letting $N_{x}=10,20,40$ and $50$. The $l_{2}$ errors $\mathcal{E}_{l_{2}}$ between numerical solutions and the reference solution for $\varepsilon=1$ are shown in Fig. 3, where the first-order convergence is clearly observed. For other $\varepsilon$, the $l_{2}$ errors $\mathcal{E}_{l_{2}}$ between the numerical solutions using $N_{x}=25,50,100$ and $200$ and the reference solution is illustrated in Fig. 4. For all these $\varepsilon$, the first-order convergence is obviously detected, indicating the efficiency of our AP-PIC scheme and a uniform first-order accuracy in space for different orders of $\varepsilon$. Therefore, our numerical scheme satisfies the AP property.
@@
Figure 3: (AP property for the RF-VPME in Sec. 5.2). The $l_{2}$ errors $\mathcal{E}_{l_{2}}$ of density $n$, velocity $u$ and potential $\phi$ for $\varepsilon=1$ at $t=0.2$. Here, the $x$-axis is $\log(\Delta x)$, and the $y$-axis is $\log(\mathcal{E}_{l_{2}})$. The mesh sizes in space are $N_{x}=10,20,40$ and $50$.
@@
Figure 4: (AP property for the RF-VPME in Sec. 5.2). The $l_{2}$ errors $\mathcal{E}_{l_{2}}$ of density $n$, velocity $u$ and potential $\phi$ for $\varepsilon=0.1,0.01$ and $10^{-4}$ at $t=0.2$. Here, the $x$-axis is $\log(\Delta x)$, and the $y$-axis is $\log(\mathcal{E}_{l_{2}})$. The mesh sizes in space are $N_{x}=25,50,100$ and $200$.
@@
In the AP-PIC method, the number of total particles is set as $N_{\rm total}=1.6\times 10^{6}$, with $N_{c}=200$ times of repetition for reduction of noise. Let the CFL number be ${\rm CFL}=0.2$. We use $N_{x}=100$ to numerically solve the Quasi-VP and RF-VPME systems. The relative errors $\mathcal{E}_{r}$ between the solutions of two systems for $\varepsilon=0.1,0.01,10^{-4}$ are shown in Fig. 5. One can observe that the smaller $\varepsilon$ is, the smaller the relative errors between the two systems are, which satisfies our expectation since the difference between the RF-VPME and Quasi-VP should be decreasing as $\varepsilon$ approaches zero. In Table 1, we show the relative errors of $n$, $u$ and $\phi$ at $t=5$ for $\varepsilon=0.1,0.01,10^{-4}$, where the same trend of decreasing errors is detected. Therefore, for long-time simulations, our scheme for solving the RF-VPME system is stable and the solution converges to that of the quasineutral limiting equation as $\varepsilon$ becomes small. In addition, from Table 1 we notice that the space discretizations do not need to resolve the Debye length as it stays fixed ($N_{x}=100)$ for different ranges of $\varepsilon$, which validates again the AP property of our numerical scheme.
@@
Figure 5: (Long time evolution of RF-VPME in Sec. 5.2). The relative errors $\mathcal{E}_{r}$ of density $n$, velocity $u$ and potential $\phi$ with $\varepsilon=0.1,0.01$ and $10^{-4}$ between RF-VPME and Quasi-VP models, from $t=0$ to $t=15$. Here, the $x$-axis is time $t$, and the $y$-axis is $\log(\mathcal{E}_{r})$. The mesh size in space is $N_{x}=100$.
@@
Table 1: (Relative errors between RF-VPME and Quasi-VP system at $t=5$ in Sec. 5.2). The relative errors of density $\mathcal{E}_{r}(n)$, velocity $\mathcal{E}_{r}(u)$ and potential $\mathcal{E}_{r}(\phi)$ for $\varepsilon=0.1,0.01,10^{-4}$ at $t=5$. The mesh size in space is $N_{x}=100$.
In the last test of this section, we investigate the $l_{2}$ errors between $n_{i}$ and $e^{\phi}$ as time evolves and for different $\varepsilon$ in the model. In the AP-PIC method, we set $N_{\rm total}=1.6\times 10^{6}$, $N_{c}=200$, $N_{x}=100$, and the CFL number is ${\rm CFL}=0.2$. Three cases of $\varepsilon=1,0.01,10^{-4}$ are considered. In each simulation, we compute $l_{2}$ errors of $(n_{i}-e^{\phi})$ and obtain the error values by taking the average of these $N_{c}$ groups of errors. In Fig. 6, we notice that as $\varepsilon$ becomes smaller, the errors between $n_{i}$ and $e^{\phi}$ decrease corresponding, especially at the longer time when errors saturate. This is because that the equation (3.11), together with initial conditions (3.12) that we used in the AP-PIC scheme, is equivalent to the original Poisson equation (3.13), which leads to the relation $n_{i}=e^{\phi}$ when $\varepsilon$ vanishes.
Figure 6: (Long time evolution for the errors between $n_{i}$ and $e^{\phi}$ in Sec. 5.2). The $l_{2}$ errors $\mathcal{E}_{l_{2}}$ of $(n_{i}-e^{\phi})$ for $\varepsilon=1,0.01,10^{-4}$ from $t=0$ to $t=15$. Here, the $x$-axis is time $t$, and the $y$-axis is $\log(\mathcal{E})$. The mesh size in space is $N_{x}=100$.
5.3 The RF-VPME system: uncertain problem
In this section, we extend to study the RF-VPME system with uncertain parameters and investigate the efficiency of our bi-fidelity (BF) approach as introduced in Section 4. The high-fidelity (HF) model is the RF-VPME system (3.15)–(3.16) solved by the AP-PIC method shown in Fig. 1. The low-fidelity (LF) model is chosen as the Euler system (4.3) computed by using the Lax-Friedrichs scheme.
To evaluate the accuracy of our method, we define errors between the BF and HF solutions at time $t$:

### v2

@@ 4. A bi-fidelity method for model with random parameters
In this section, we briefly review the bi-fidelity method. Assume the expensive high-fidelity model and the cheap low-fidelity model are available to generate the high-fidelity solution $\mathbf{U}^{H}(z)$ and the low-fidelity solution $\mathbf{U}^{L}(z)$ respectively, for any given parameter $z$. We refer to [45, 55] for more details and summarize the bi-fidelity method as follows: {lgrthm} A bi-fidelity approximation
0:
@@
1: Select a sample set $\Gamma=\{z_{1},z_{2},\ldots,z_{M}\}\subset I_{z}$, and run the low-fidelity model $\mathbf{U}^{L}(z_{j})$ for each $z_{j}\in\Gamma$.
@@
2: Select $r$ “important” points from $\Gamma$ and denote it by $\gamma=\{z_{1},\cdots z_{r}\}\subset\Gamma$. Construct the low-fidelity approximation space $\mathscr{U}^{L}(\gamma)$.
@@
3: Run high-fidelity simulation at each point in the selected set $\gamma$. Construct the high-fidelity approximation space $\mathscr{U}^{H}(\gamma)$.
3:
4: For any $z$, compute the low-fidelity solution $U^{L}(z)$ and the corresponding low-fidelity coefficients:
$\mathbf{U}^{L}(z)\approx\mathcal{P}_{\mathscr{U}^{L}\left(\gamma\right)}\left[\mathbf{U}^{L}(z)\right]=\sum_{k=1}^{r}c_{k}^{L}(z)\mathbf{U}^{L}\left(z_{k}\right).$ (4.1)
5: Construct the bi-fidelity approximation by applying the same approximation rule as in the low-fidelity model:
$\mathbf{U}^{B}(z)=\sum_{k=1}^{r}c_{k}^{L}(z)\mathbf{U}^{H}(z_{k}).$ (4.2)
@@
Our sought bi-fidelity approximation of $u^{H}$ is constructed by equation (4.2). In the offline stage, one employs the cheap low-fidelity model to select the most important parameter points $\gamma=\{z_{1},\cdots,z_{r}\}$ by the greedy procedure [18]. In the online stage, for any sample point $z\in I_{z}$, we project the low-fidelity solution $\mathbf{U}^{L}(z)$ onto the low-fidelity approximation space $\mathscr{U}^{L}(\gamma)$ by using (4.1), where $\mathcal{P}_{\mathscr{U}^{L}(\gamma)}$ is the projection operator onto $\mathscr{U}^{L}(\gamma)$ with coefficients $c_{k}$ computed by the Galerkin approach:
${\bf G}^{L}{\bf c}={\bf f},\,\text{\penalty 10000\ with\penalty 10000\ }{\bf f}=(f_{k})_{1\leq k\leq r}\,\text{\penalty 10000\ and\penalty 10000\ }f_{k}=\langle\mathbf{U}^{L}(z),\mathbf{U}^{L}(z_{k})\rangle^{L},$
where ${\bf G}^{L}$ is the Gramian matrix for $\mathbf{U}^{L}(\gamma)$.
@@
High-fidelty solver: We desire to solve the VPME system with multi-dimensional uncertain parameters, thus in our bi-fidelity method, the VPME system (2.1c) is considered as the high-fidelity model. We employ the AP-PIC method studied in Subsection 3.3 as our high-fidelity solver.
@@
Low-fidelity solver: We choose the macroscopic Euler system (2.4) that characterizes a simpler physics as our low-fidelity model in the bi-fidelity method. We use the finite difference method (FDM) to discretize the spatial space and forward Euler scheme for the time discretization. Consider the conservation form of (2.4) for one-dimensional space and velocity:
$\frac{\partial\mathbf{U}}{\partial t}+\frac{\partial\mathbf{F}(\mathbf{U})}{\partial x}=0,\quad\mathbf{U}=\begin{bmatrix}n\\ nu\end{bmatrix},\quad\mathbf{F}(\mathbf{U})=\begin{bmatrix}nu\\ nu^{2}+n\end{bmatrix}$. (4.3)
@@
where n=n(x,t),u=u(x,t)n=n(x,t),\>u=u(x,t) are the ion density and velocity given in (2.3) in the computational domain $\left[x_{L},x_{R}\right]$. Let $h$ be a uniform space step and $\tau$ be a uniform time step, then $x_{j}=jh,\>j=0,1,2,\dots,N_{l}$ with $x_{L}=x_{0},\>x_{R}=x_{N_{l}}$ and $t^{n}=n\tau,\>n=1,2,\dots,M$.
@@
In our flux construction, Lax-Friedrichs scheme–as one of the most popular finite difference methods for solving hyperbolic partial differential equations [38, 52, 53] is used. Let $\mathbf{U}^{n}_{j}=\mathbf{U}(x_{j},t^{n})$ be the approximated values of $\mathbf{U}$ at position $x_{j}$ and time $t^{n}$. The time discretization for (4.3) becomes
$\mathbf{U}^{n+1}_{j}=\mathbf{U}^{n}_{j}-\frac{\tau}{h}(\mathbf{F}^{n}_{j+\frac{1}{2}}-\mathbf{F}^{n}_{j-\frac{1}{2}}),$ (4.4)
where $\mathbf{F}^{n}_{j+\frac{1}{2}}$ is given by
$\mathbf{F}^{n}_{j+\frac{1}{2}}=\frac{1}{2}\left(\mathbf{F}(\mathbf{U}^{n}_{j+1}\right)+\mathbf{F}(\mathbf{U}^{n}_{j}))-\frac{h}{2\tau}(\mathbf{U}^{n}_{j+1}-\mathbf{U}^{n}_{j}).$ (4.5)
@@
$\mathbf{U}^{n+1}_{j}=\frac{\mathbf{U}^{n}_{j+1}+\mathbf{U}^{n}_{j-1}}{2}-\frac{\tau}{2h}(\mathbf{F}(\mathbf{U}^{n}_{j+1})-\mathbf{F}(\mathbf{U}^{n}_{j-1})).$ (4.6)
@@
5. Numerical examples
In our numerical experiments, we first study the deterministic VPME system, then extend to the problems with random uncertainties and validate the efficiency of our bi-fidelity method. For all the following experiments, the regularization parameter $\eta$ in (3.1d) and the error tolerance $\eta_{1}$ in the Newton’s method (3.30) for solving the reformulated Poisson are set as $\eta=1,\,\eta_{1}=1\times 10^{-8}$. We denote the total number of particles for a single PIC run as $N_{\text{tol}}$, and the notation $N_{c}$ represents the number of independent numerical simulations that are performed under the fixed setting and numerical parameters. The purpose of running the same test $N_{c}$ times is to reduce the statistical noise inherent in the PIC method through the ensemble averaging technique. For notational convenience, we name the reformulated VPME system (3.15)–(3.16) as the RF-VPME system, and its quasineutral limit (3.17)–(3.18) as the Quasi-VP system.
5.1. The RF-VPME system: an explicit scheme
@@
Set the final time as $t=2$ and $\varepsilon=0.01$ in the model. Let $v\in[-4,4]$ with uniform mesh and the number of grid points be $N_{v}=800$, the spatial mesh size be $N_{x}=200$ and time step be $\Delta t=0.2\Delta x$. The density $n$ and velocity $u$ are shown in Figure 2, where the blue line represents solution of the explicit scheme (3.36) based on building a penalty for the nonlinear term $e^{\phi^{m+1}}$, and the red line stands for the implicit scheme using Newton’s iteration (3.29). The solutions of both methods match well, and both schemes are stable for long time. In our future work, we will incorporate this explicit scheme with the PIC method for solving the Vlasov equation (2.1a) and develop higher-order AP schemes.
Figure 2. (Two schemes for solving the reformulated Poisson, $\varepsilon=0.01$). Density $n$ (left) and velocity $u$ (right) at $t=2$ by an implicit scheme (3.23) with Newton’s iteration (red) and an explicit scheme based on penalty (3.36) (blue).
5.2. The RF-VPME system: deterministic problem
@@
We particularly study the cases when $\varepsilon=1,0.1,0.01$ and $10^{-4}$. The final time is $t=0.2$. When $\varepsilon=1$, the numerical solution is obtained by letting $N_{x}=10,20,40$ and $50$. The $l_{2}$ errors $\mathcal{E}_{l_{2}}$ between numerical solutions and the reference solution for $\varepsilon=1$ are shown in Figure 3, where the first-order convergence is clearly observed. For other $\varepsilon$, the $l_{2}$ errors $\mathcal{E}_{l_{2}}$ between the numerical solutions using $N_{x}=25,50,100$ and $200$ and the reference solution is illustrated in Figure 4. For all these $\varepsilon$, the first-order convergence is obviously detected, indicating the efficiency of our AP-PIC scheme and a uniform first-order accuracy in space for different orders of $\varepsilon$. Therefore, our numerical scheme satisfies the AP property.
@@
Figure 3. (AP property for the RF-VPME in Subsection 5.2). The $l_{2}$ errors $\mathcal{E}_{l_{2}}$ of density $n$, velocity $u$ and potential $\phi$ for $\varepsilon=1$ at $t=0.2$. Here, the $x$-axis is $\log(\Delta x)$, and the $y$-axis is $\log(\mathcal{E}_{l_{2}})$. The mesh sizes in space are $N_{x}=10,20,40$ and $50$.
@@
Figure 4. (AP property for the RF-VPME in Subsection 5.2). The $l_{2}$ errors $\mathcal{E}_{l_{2}}$ of density $n$, velocity $u$ and potential $\phi$ for $\varepsilon=0.1,0.01$ and $10^{-4}$ at $t=0.2$. Here, the $x$-axis is $\log(\Delta x)$, and the $y$-axis is $\log(\mathcal{E}_{l_{2}})$. The mesh sizes in space are $N_{x}=25,50,100$ and $200$.
@@
In the AP-PIC method, the number of total particles is set as $N_{\rm total}=1.6\times 10^{6}$, with $N_{c}=200$ times of repetition for reduction of noise. Let the CFL number be ${\rm CFL}=0.2$. We use $N_{x}=100$ to numerically solve the Quasi-VP and RF-VPME systems. The relative errors $\mathcal{E}_{r}$ between the solutions of two systems for $\varepsilon=0.1,0.01,10^{-4}$ are shown in Figure 5. One can observe that the smaller $\varepsilon$ is, the smaller the relative errors between the two systems are, which satisfies our expectation since the difference between the RF-VPME and Quasi-VP should be decreasing as $\varepsilon$ approaches zero. In Table 1, we show the relative errors of $n$, $u$ and $\phi$ at $t=5$ for $\varepsilon=0.1,0.01,10^{-4}$, where the same trend of decreasing errors is detected. Therefore, for long-time simulations, our scheme for solving the RF-VPME system is stable and the solution converges to that of the quasineutral limiting equation as $\varepsilon$ becomes small. In addition, from Table 1 we notice that the space discretizations do not need to resolve the Debye length as it stays fixed ($N_{x}=100)$ for different ranges of $\varepsilon$, which validates again the AP property of our numerical scheme.
@@
Figure 5. (Long time evolution of RF-VPME in Subsection 5.2). The relative errors $\mathcal{E}_{r}$ of density $n$, velocity $u$ and potential $\phi$ with $\varepsilon=0.1,0.01$ and $10^{-4}$ between RF-VPME and Quasi-VP models, from $t=0$ to $t=15$. Here, the $x$-axis is time $t$, and the $y$-axis is $\log(\mathcal{E}_{r})$. The mesh size in space is $N_{x}=100$.
@@
Table 1. (Relative errors between RF-VPME and Quasi-VP system at $t=5$ in Subsection 5.2). The relative errors of density $\mathcal{E}_{r}(n)$, velocity $\mathcal{E}_{r}(u)$ and potential $\mathcal{E}_{r}(\phi)$ for $\varepsilon=0.1,0.01,10^{-4}$ at $t=5$. The mesh size in space is $N_{x}=100$.
In the last test of this section, we investigate the $l_{2}$ errors between $n_{i}$ and $e^{\phi}$ as time evolves and for different $\varepsilon$ in the model. In the AP-PIC method, we set $N_{\rm total}=1.6\times 10^{6}$, $N_{c}=200$, $N_{x}=100$, and the CFL number is ${\rm CFL}=0.2$. Three cases of $\varepsilon=1,0.01,10^{-4}$ are considered. In each simulation, we compute $l_{2}$ errors of $(n_{i}-e^{\phi})$ and obtain the error values by taking the average of these $N_{c}$ groups of errors. In Figure 6, we notice that as $\varepsilon$ becomes smaller, the errors between $n_{i}$ and $e^{\phi}$ decrease corresponding, especially at the longer time when errors saturate. This is because that the equation (3.11), together with initial conditions (3.12) that we used in the AP-PIC scheme, is equivalent to the original Poisson equation (3.13), which leads to the relation $n_{i}=e^{\phi}$ when $\varepsilon$ vanishes.
Figure 6. (Long time evolution for the errors between $n_{i}$ and $e^{\phi}$ in Subsection 5.2). The $l_{2}$ errors $\mathcal{E}_{l_{2}}$ of $(n_{i}-e^{\phi})$ for $\varepsilon=1,0.01,10^{-4}$ from $t=0$ to $t=15$. Here, the $x$-axis is time $t$, and the $y$-axis is $\log(\mathcal{E})$. The mesh size in space is $N_{x}=100$.
5.3. The RF-VPME system: uncertain problem
In this section, we extend to study the RF-VPME system with uncertain parameters and investigate the efficiency of our bi-fidelity (BF) approach as introduced in Subsection 4. The high-fidelity (HF) model is the RF-VPME system (3.15)–(3.16) solved by the AP-PIC method shown in Figure 1. The low-fidelity (LF) model is chosen as the Euler system (4.3) computed by using the Lax-Friedrichs scheme. To evaluate the accuracy of our method, we define errors between the BF and HF solutions at time $t$: